# Turbulence-induced anti-Stokes flow: experiments and theory

### v1

@@ 3.1 Flow and wave characteristics
When comparing turbulent quantities for the various cases one should bear in mind that these are measured at a fixed position in space, whereas the turbulence becomes gradually weaker as it travels downstream because of dissipation. The change in current is an integrated effect of wave-turbulence interactions that occurred upstream, where the turbulence intensity is in general a little higher. This is true of all cases, but because of the lower mean velocity, the turbulence that reaches the field of view in cases 3.A and 3.B has decayed for longer (a detailed study of turbulent decay in our lab was reported by Jooss et al., 2021).
In previous experiments by Nepf & Monismith (1991) and Klopman (1994), secondary motion in the form of a pair of streamwise rolls were measured when waves were propagated on a current, triggered by the CL2 mechanism. We could measure the cross-plane flow in Experiment 1 and see traces of what could be such a flow, with mean spanwise velocities of $1$ cm or less. However, since we have a wider channel, shorter test section and faster flow than either of these, water from the sidewall boundary layers only reach a little way into the main channel, leaving a channel at least $80$ cm wide at the centre of the flow undisturbed at the point of measurement, even wider further upstream. We conclude that this phenomenon has negligible effect on our results.
Other studies in the same facility (without waves) give us confidence that the waves are essentially unaffected by turbulence created in the bottom and wall boundary layers (Jooss et al., 2021). In fact, the recent study of Asadi et al. (2025) in the same facility showed that the addition of active-grid turbulence acts to contain the boundary-layer turbulence closer to the walls.
@@
Due to differences in the way the experimental data were acquired, no single method for estimating the integral scale $L_{x}^{x}$ can be applied to all cases. Estimating the integral lengthscales from experimental data uniquely and quantitatively is notoriously difficult, and the various methods in common use produce quantitatively different results. Additional challenges pertain to active-grid turbulence due to the slow spatial decay of the autocorrelation function (Puga & LaRue, 2017; Mora et al., 2019). In Experiment 1, the integral scale was estimated with a zero-crossing method as described by Mora & Obligado (2020), and we used the same method to calculate the integral scale for Experiment 3, as listed in table 3. Since the data in Experiment 2 are not time resolved, the same method cannot be applied there, and using another, spatially based method would not give directly comparable numbers, so we provide no integral scale for Experiment 2. We note that integral scales are significantly shorter in Experiment 3 than in Experiment 1 at similar turbulence levels, which can be explained by the lower mean flow velocity $U_{0}$.
@@ 3.1.1 Wave group measurements (Experiment 1)
The mean flow and turbulence statistics without waves were evaluated over the first interval, where any influence from waves can be assumed to be negligible. The vertical profile of the mean flow was approximately constant in the spanwise and vertical directions over the field of view with the absolute value of the mean flow (positive in the negative $x$-direction), $U_{0}$, found from averaging as described above.
The characteristic peak wave amplitude $a_{\mathrm{p}}$ and the group temporal width $\tau$ were estimated from the ensemble-averaged amplitude envelope of the wave groups as measured by the probes near the SPIV/LIF laser sheet as seen in figure 2b). The average envelope was fitted to a Gaussian function of the form
@@
with $t_{\mathrm{p}}$ the temporal location of the group peak. The peak wave steepness $\epsilon_{0\mathrm{p}}=a_{\mathrm{p}}k_{0}$.
We define an intrinsic temporal group width $\tau_{0}$, listed in table 3, defined as:
@@
We will find it useful to consider an “effective interaction time”, the duration for which the turbulent current has been influenced by waves before it reaches the point where it is measured, taking into account that the wave amplitudes vary throughout the group. Under the assumption that the instantaneous influence of waves on passing turbulence is proportional to $a(t)^{2}$, a reasonable estimate for the cases 1.A1.D is
$T_{\mathrm{int}}=\frac{1}{a_{\mathrm{p}}^{2}}\int_{-\infty}^{\infty}\tilde{a}(t)^{2}\mathrm{d}t=\sqrt{\pi}\tau_{0},$ (6)
where $\tilde{a}(t)$ is the time-dependent amplitude seen by the turbulence, i.e., with $\tau_{0}$ replacing $\tau$ in (4). The interaction times, given in table 3, are in the vicinity of $7$ times the intrinsic period of the carrier wave, $\varpi_{0}=2\pi/(k_{0}c(k_{0}))$.
@@ 3.1.2 Measurements with regular waves (Experiments 2 and 3)
The wave steepness $\epsilon_{0}$ was calculated using the average wave amplitude from the wave probe measurements in the proximity of the PIV measurement location. The amplitude of each individual wave oscillation varied slightly during the experiments, especially in cases with the highest level of turbulence, likely due to wave-turbulence interactions (e.g., Smeltzer et al., 2023). The variation is not of central interest to the present study, and thus only the mean steepness value is reported.
The regular waves were present throughout the entire test section during the experiments. The grid-generated and measured Turbulence thus interacted with the waves over a length $L_{\mathrm{FOV}}$ as given in Sections 2.1.2 and 2.1.3, with associated interaction time $T_{\mathrm{int}}=L_{\mathrm{FOV}}/U_{0}$.
@@ 3.2 Measured change in Eulerian-mean velocity
The measured changes in Eulerian-mean velocities presented below are in the order of millimetres per second, yet, while this is the same order of magnitude as our PIV measurement accuracy, one should bear in mind that the variance of an average from hundreds of independent measurements is far lower than that of single measurements. It is crucial that a careful analysis of errors and statistical convergence be performed in order to establish confidence that our results are accurate and reliable. In appendix B, we report results of these tests, supported by further data in the Supplementary Materials.
@@ 3.2.1 Velocity change after the passage of wave groups (Experiment 1)
Figure 3: Change in Eulerian-mean current due to the passage of a wave group. The waves travelled in the positive $x$-direction, against the current. (a) An example of mean streamwise velocity depth profile $U_{I}(z)$ for intervals $I=\{1,2,3\}$, here for case 1.D; (b) Mean streamwise velocity difference $\Delta U=U_{3}-U_{1}$ as a function of depth for flow cases 1.A1.D as denoted in the legend. Error bars are omitted for visibility — see analysis in Appendix B; (c) Slope relative to the Stokes drift gradient (a prime denotes derivation with respect to $z$). Light smoothing (moving average with window size $8$ mm) was applied to the curves for better visibility.
@@
We now consider the measured Eulerian-mean velocity change for Experiment 1, i.e., cases 1.A1.D. Vertically sheared flows in our laboratory have been found to be stable and therefore the wave-modified Eulerian-mean velocity profile remains close to unchanged throughout interval 3 (also found to be true for strongly sheared currents, see Section 4 of Pizzo et al., 2023). In figure 3(a) we show as an example the mean streamwise velocity profiles for case 1.D during the three measurement intervals, denoted $U_{I}(z)$ for intervals $I=\{1,2,3\}$ (see figure 2(b)), respectively. As can be seen, the mean flow speed had increased in the direction opposite to wave propagation after the passage of the wave groups. Similar plots for all cases are provided in the Supplementary Materials. Error bars are omitted from the figure for reasons of visibility, but discussed in Appendix B. The net Eulerian-mean flow change $\Delta U=U_{3}-U_{1}$ is shown in figure 3(b) for flow cases 1.A1.D as expressed in the legend. For all cases, $U_{3}-U_{1}<0$ near the surface, and decays to small absolute values at depths $|k_{0}z|\gtrsim 1$.
@@
Interestingly, $\Delta U$ clearly depends on the level of ambient turbulence; while the waves have approximately equal properties in Cases 1.A, 1.B, 1.C.1 and 1.D, the velocity change is far higher at the highest turbulence level (case 1.D) than at the lowest level (case 1.A), with the intermediate cases 1.B and 1.C.1 in between. Moreover, comparing cases 1.C.1 and 1.C.2 where two wave steepnesses are tested on the same turbulent current indicates a positive correlation between $\epsilon_{0}$ and current change $\Delta U$. Put together, the results in figure 3(b) provide a strong indication that the change in current is due to an interaction between waves and turbulence.
@@
We can exclude the possibility that the measured $\Delta U$ is due to the Eulerian return flow under groups of waves, as measured by van den Bremer et al. (2019). The return flow follows the group, i.e., it is very weak in intervals 2 and 3 which lie outside the main group itself. It is also uniform in depth, whereas the current measured in figure 3a is strongly depth dependent.
@@
It is instructive to plot the ratio between $\mathrm{d}(\Delta U)/\mathrm{d}z$ and $-\mathrm{d}u_{s}/\mathrm{d}z$ as a function of $z$, shown in figure 3(c), because it gives some indication of how far the turbulent flow has transitioned towards a new, quasi-equilibrium state. We will later present theory and evidence that at the end of the ‘spin-up’ period, this ratio should, in the final state, be approximately equal to $u_{\infty}^{2}/w_{\infty}^{2}$ nearest the surface. Since our bulk turbulence is slightly anisotropic (in the cases reported in Jooss et al. (2021), $1.2\lesssim u_{\infty}^{2}/w_{\infty}^{2}\lesssim 1.4$), a final current change $|\Delta U|\gtrsim|u_{s}|$ is expected nearest the surface. We shall later see that in the regular-wave cases where the equilibrium is likely reached, this relation holds well. Since none of the changes in currents in cases 1.A1.D are close to reaching these values, it seems that the passing of the wave group has not led to a wave-turbulence interaction of sufficient duration for a final state to be reached, and the flow is still relatively early in the ‘spin-up’ stage.
@@
There are at least two striking observations to make in figure 3(c). First, with the exception of the low-turbulence case, 1.A, the ratio between the slopes is close to constant with depth, which illustrates that $\Delta U\sim\exp(2k_{0}z)$ near the surface in these cases. The scaling is not perfect, particularly at the shallowest depths; this should not be surprising since the depth dependence of wave-turbulence interaction should scale not only with the wavelength, but also the turbulent integral scale, which delimits the vertical extent of the topmost layer where the kinematic boundary condition at the surface begins to limit the vertical extent of turbulent eddies (the blocking effect, see, e.g., Teixeira & Belcher, 2002). Second, while the value of $\Delta U(z)$ after the passing of a group depends strongly on the turbulence level and steepness, there is no such trend for the relative slope ($-\Delta U^{\prime}(z)/u_{s}^{\prime}(z)$), Case 1.A excepted.
@@
In Section 4.4 we will develop a RDT model describing the early onset of wave-turbulence interaction, which we can compare to the measurements in figure 3b, given this evidence that the combined wave/turbulence flow is still far from fully developed in Experiment 1.
@@
In conclusion, the evidence suggests that the change in Eulerian-mean current observed in our experiments is due to wave-turbulence interaction, and increases with increasing turbulence and increasing steepness.
@@ 3.3 Velocity change under regular waves (Experiments 2 and 3)
Cases 2.A3.B all consider turbulence interacting with regular (i.e., continuous and periodic) waves. For cases 2.A1–2.B2, we evaluate the mean streamwise velocity in the presence of waves, and subtract off the mean velocity profile from the ambient flow case without waves. This velocity difference we define as $\Delta U(z)$. Although the flow is wavy, the time series is long enough for the Eulerian-mean current, averaged over all $2000$ PIV images, to be well converged in the sense that further measurements would affect it insignificantly. See details in Appendix B.
@@
The turbulent current is affected by the waves during the time it takes it to traverse the test section, a distance $L_{\mathrm{FOV}}$ as defined in Sections 2.1.2 and 2.1.3. The interaction time is thus $T_{\mathrm{int}}=L_{\mathrm{FOV}}/U_{0}$, listed in table 3. Due to the slower flow speed. cases 3.A and 3.B interact for considerably longer than 2.A and 2.B, both in terms of absolute time (in seconds) and in terms of number of intrinsic wave periods ($2\pi/\varpi_{0}$).
@@
Figure 4: The wave-induced current $\Delta U$ under regular waves as a function of depth $k_{0}z$ for cases 2.A1, 2.A2, 2.B1, 2.B2, 3.A and 3.B in panels (a)–(f), respectively. For each case, different wave steepness values are shown as indicated in the legend. The dashed lines are the theoretical Stokes drift profiles at the same location for each case, shown as $-u_{s}(z)$, that is, with opposite sign to the Stokes drift. The filled circles at $k_{0}z=-4$ and 00 indicate the theoretical value of the Eulerian return flow, $u_{\mathrm{rf}}$.
@@
In figure 4 we show $\Delta U(z)$ for cases 2.A3.B, where the different wave steepness values $\epsilon_{0}$ are labelled in the legend. Several characteristic behaviours can be observed. First, all graphs have similar dependence on depth, with the highest absolute value nearest to the surface, decreasing down to a level where $k_{0}z$ is roughly in the range between $-1$ and $-2$, then turning and slowly becoming more negative again. (The non-monotonicity (turning) would not be visible in figure 3, where measurements were only made for $k_{0}z>-1.2$). Indeed, some cases see the induced Eulerian-mean current take positive values over a certain depth range. (One may note that the current profile is similar in shape and magnitude to those of Rashidi et al., 1992). Second, there is again a clear tendency that higher steepness leads to a larger vertical variation of $\Delta U(z)$. Finally, there appears to be a near-constant offset between graphs (an addition to the mean flow), which increases with steepness, and we will find in Section 4.1 that it can be explained, at least in part, as the approximately depth-uniform Eulerian return flow $u_{\mathrm{rf}}$.
@@
The reverse of the theoretical Stokes drift $-u_{s}(z)$ from (1) is plotted in all panels of figure 4 with dashed lines of the same colour for comparison, calculated for each case. For all cases we note that $|\Delta U|<u_{s}$ nearest the surface, but that the depth variation $\mathrm{d}|\Delta U|/\mathrm{d}z\sim\mathrm{d}u_{s}/\mathrm{d}z$, is higher in some cases, smaller in others. We will return to these points later when comparing the results to theory.
@@
To further illustrate how the Eulerian-mean-current change depends on wave steepness, we plot the value of $\Delta U$ at a set reference depth as a function of steepness $\epsilon_{0}=k_{0}a_{0}$ in figure 5. The value at the shallowest depth available for all cases is used, i.e. $k_{0}z=-0.27$. For cases 2.A.1 and 2.B.1, $\Delta U$ appears to scale as $\epsilon_{0}^{2}$, indicated by the dashed line. Cases 2.A2 and 2.B2 do not adhere to the scaling, possibly due to the high wave steepness values involved, so that interactions of order $\epsilon_{0}^{3}$ and higher may become significant. A further possible explanation is the lack of scale separation: the wavelength is likely only slightly larger than the integral scale in these two cases (based on comparison with cases 1.A1.D in table 3, which have similar $U_{0}$ — unfortunately, a direct comparison of $L_{x}^{x}$ is not possible, as explained). The comparatively large integral scale and high turbulence levels mean that stronger angular scattering of waves on turbulent velocity changes is to be expected (Villas Bôas & Young, 2020; Smeltzer et al., 2023), a process which does not scale with steepness. Figure 5 should be interpreted only qualitatively, since the value of $\Delta U$ at a constant value of $k_{0}z$ is not entirely comparable between cases with different turbulence properties. As discussed in connection with figure 3(c) and at length in Section 4.4, $\Delta U$ depends not only on $k_{0}$ but also on the turbulent integral scale and anisotropy, which varies between cases.
@@
Figure 5: The wave-induced current under regular waves at the ‘reference’ depth $k_{0}z=-0.27$ as a function of wave steepness for cases 2.A3.B as indicated in the legend. The dashed line is proportional to $\epsilon_{0}^{2}$.

### v2

@@ 3.1 Flow and wave characteristics
When comparing turbulent quantities for the various cases one should bear in mind that these are measured at a fixed position in space, whereas the turbulence becomes gradually weaker as it travels downstream because of dissipation. The change in Eulerian-mean current is an integrated effect of wave–turbulence interactions that occurred upstream, where the turbulence intensity is in general a little higher. This is true of all cases, but because of the lower mean velocity, the turbulence that reaches the field of view in cases 3.A and 3.B has decayed for longer. A detailed study of turbulent decay in our lab with similar flow conditions as Experiment 1 was reported by Jooss et al. (2021). They found that for all cases, once the turbulence is fully developed TKE decays approximately as $\sim 1/x$ downstream for all cases. Thus, although the absolute values of TKE and Reynolds stresses are different at different positions, their relative magnitudes are expected to be constant. We consider only trends with respect to TKE, and ratios between Reynolds stresses which it seems reasonable to assume are well captured.
In previous experiments by Nepf & Monismith (1991) and Klopman (1994, 1997), secondary motion in the form of a pair of streamwise rolls were measured when waves were propagated on a current, triggered by the CL2 mechanism. We could measure the cross-plane flow in Experiment 1, and while some weak in-plane mean flow was observed, of the order of $1$ cm/s or less, the pattern was not consistent with a coherent pair of rolls. The resulting convection, in magnitude and direction, would be insufficient to convect boundary-layer turbulence into the near-surface regions where turbulence and waves interact to any noticeable degree. During the travel from grid to measurement position, the vortex force near the boundaries has far less time to accelerate a similar amount of water as in Klopman (1994) and would likely not have time to develop. It is not perhaps so surprising that secondary flows would be different given that geometrical size and aspect ratios are rather different, and the current’s travel time from grid to measurement point is less than half that in either of the mentioned flows even for our slowest flow in Experiment 3.
@@
It was found by Smeltzer et al. (2023) that the turbulent integral length scale, characteristic of the largest turbulent structures, is important for wave–turbulence interactions, especially so for the scatting of waves by turbulence. An advantage of our active-grid turbulence generation is that the integral scale can be varied independently of TKE. The most pertinent integral scale is related to correlation of streamwise turbulent velocity for points separated in the streamwise direction, $L_{x}^{x}$. Due to differences in the way the experimental data were acquired, no single method for estimating $L_{x}^{x}$ can be applied to all cases. Estimating the integral lengthscales from experimental data uniquely and quantitatively is notoriously difficult, and the various methods in common use produce quantitatively different results. Additional challenges pertain to active-grid turbulence due to the slow spatial decay of the autocorrelation function (Puga & LaRue, 2017; Mora et al., 2019). In Experiment 1, the integral scale was estimated with a zero-crossing method as described by Mora & Obligado (2020), and we used the same method to calculate the integral scale for Experiment 3, as listed in table 2. Since the data in Experiment 2 have low time resolution the same method cannot be applied there, and using another, spatially based method would not give directly comparable numbers, so we provide no integral scale for Experiment 2. We note that integral scales are significantly shorter in Experiment 3 than in Experiment 1 at similar turbulence levels, which can be explained by the lower mean-flow velocity $U_{0}$.
@@ 3.1.1 Wave group measurements (Experiment 1)
The mean flow and turbulence statistics without waves presented in Table 2 were evaluated over the first interval, where any influence from waves can be assumed to be negligible. The mean flow in the negative $x$-direction had approximately constant (absolute) value $U_{0}$, found from averaging as described above. A small spanwise and vertical mean velocity was found, everywhere below $1$ cm/s, mostly much less, which we attribute to the channel flow conditioning not being perfect (achieving a perfectly straight and uniform flow in a channel of our size is very challenging). Notably, unlike the streamwise velocity, the cross-plane velocities did not change after the passage of the wave group, which implies that there is negligible contribution secondary flow due to wave–current interactions near the side-walls mentioned above.
The characteristic peak wave amplitude $a_{\mathrm{p}}$ and the observed group temporal width $\tau$ were estimated from the ensemble-averaged amplitude envelope of the wave groups as measured by the probes near the SPIV/LIF laser sheet as seen in figure 2(b). The average envelope was fitted to a Gaussian function of the form
@@
with $t_{\mathrm{p}}$ the temporal location of the group peak. The peak wave steepness is $k_{0}a_{\mathrm{p}}$.
We define an intrinsic temporal group width $\tau_{0}$, listed in table 3 as:
@@ 3.1.2 Measurements with regular waves (Experiments 2 and 3)
The wave steepness $ak_{0}$ was calculated using the average wave amplitude from the wave probe measurements in the proximity of the PIV measurement location. The amplitude of each individual wave oscillation varied slightly during the experiments, especially in cases with the highest level of turbulence, likely due to wave–turbulence interactions (e.g., Smeltzer et al., 2023). The variation is not of central interest to the present study, and thus only the mean steepness value is reported.
The regular waves were present throughout the entire test section during the experiments. The grid-generated and measured turbulence thus interacted with the waves over a length $L_{\mathrm{FOV}}$ as given in Sections 2.1.2 and 2.1.3, with associated interaction time $T_{\mathrm{int}}=L_{\mathrm{FOV}}/U_{0}$. The frequency at which turbulence encounters waves, i.e., as seen by the moving flow, equals the intrinsic wave frequency (in Hz)
$f_{\mathrm{enc}}=f_{0}+U_{0}/\lambda_{0}=\sqrt{gk_{0}}/2\pi.$ (6)
@@ 3.2 Measured change in Eulerian-mean velocity
The measured changes in Eulerian-mean velocities presented below are of the order of millimetres per second, yet, while this is the same order of magnitude as our PIV measurement accuracy, one should bear in mind that the variance of an average from hundreds of independent measurements is far lower than that of single measurements. It is crucial that a careful analysis of errors and statistical convergence be performed in order to establish confidence that our results are accurate and reliable. In appendix B, we report results of these tests, supported by further data in the Supplementary Materials.
@@ 3.2.1 Velocity change after the passage of wave groups (Experiment 1)
Figure 3: Change in Eulerian-mean current due to the passage of a wave group. The waves travelled in the positive $x$-direction, against the current. (a) An example of mean streamwise velocity depth profile before the arrival of a wave group, $U_{1}(z)$, and after the group has passed, $U_{3}(z)$, here for case 1.D; (b) mean streamwise velocity difference $\Delta U=U_{3}-U_{1}$ as a function of depth for the flow cases of Experiment 1. Error bars are omitted for visibility – see analysis in Appendix B; (c) the slope of $\Delta U(z)$ relative to the Stokes drift gradient (a prime denotes derivation with respect to $z$). Light smoothing (moving average with window size 88 mm) was applied to the curves in panel c for better visibility.
@@
We now consider the measured Eulerian-mean velocity change for Experiment 1, i.e., cases 1.A1.D. Vertically sheared flows in our laboratory have been found to be stable and therefore the wave-modified Eulerian-mean velocity profile remains close to unchanged throughout interval 3 (also found to be true for strongly sheared currents, see Section 4 of Pizzo et al., 2023). In figure 3(a) we show as an example the mean streamwise velocity profiles for case 1.D during the three measurement intervals, denoted $U_{I}(z)$ for intervals $I=\{1,2,3\}$ (see figure 2(b)), respectively. As can be seen, the mean flow speed had increased in the direction opposite to wave propagation after the passage of the wave groups. Similar plots for all cases are provided in the Supplementary Materials. Error bars are omitted from the figure for reasons of visibility, but discussed in Appendix B. The net Eulerian-mean flow change $\Delta U=U_{3}-U_{1}$ is shown in figure 3(b) for flow cases 1.A1.D as expressed in the legend. For all cases, $\Delta U<0$ near the surface, and decays to small absolute values at depths $|k_{0}z|\gtrsim 1$.
@@
Interestingly, $\Delta U$ clearly depends on the level of ambient turbulence; while the waves have approximately equal properties in Cases 1.A, 1.B, 1.C.1 and 1.D, the velocity change is far higher at the highest turbulence level (case 1.D) than at the lowest level (case 1.A), with the intermediate cases 1.B and 1.C.1 in between. Moreover, comparing cases 1.C.1 and 1.C.2 where two wave steepnesses are tested on the same turbulent current indicates a positive correlation between $k_{0}a_{\mathrm{p}}$ and current change $\Delta U$. Put together, the results in figure 3(b) provide a strong indication that the change in current is due to an interaction between waves and turbulence.
@@
We can exclude the possibility that the measured $\Delta U$ is due to the Eulerian return flow under groups of waves, as measured by van den Bremer et al. (2019). The return flow follows the group, i.e., it is very weak in intervals 2 and 3 which lie outside the main group itself. It is also near-uniform in depth, whereas the current measured in figure 3a is strongly depth dependent.
@@
It is instructive to plot the ratio between $\mathrm{d}(\Delta U)/\mathrm{d}z$ and $\mathrm{d}u_{s}/\mathrm{d}z$ as a function of $z$, shown in figure 3(c), because it gives some indication of how far the turbulent flow has transitioned towards a new, quasi-equilibrium state. We will later present theory and evidence that at the end of the ‘spin-up’ period, this ratio should, in the final state, be approximately equal to $u_{\mathrm{rms}}^{2}/w_{\mathrm{rms}}^{2}$ nearest the surface. Since our bulk turbulence is slightly anisotropic (in the cases reported in Jooss et al. (2021), $1.2\lesssim u_{\mathrm{rms}}^{2}/w_{\mathrm{rms}}^{2}\lesssim 1.4$), a final current change $|\Delta U|\gtrsim|u_{s}|$ is expected nearest the surface. We shall later see that in the regular-wave cases where the equilibrium is likely reached, this relation holds well. Since none of the changes in currents in cases 1.A1.D are close to reaching these values, it seems that the passing of the wave group has not led to a wave–turbulence interaction of sufficient duration for a final state to be reached, and the flow is still relatively early in the ‘spin-up’ stage.
@@
There are at least two striking observations to make in figure 3(c). First, with the exception of the low-turbulence case, 1.A, the ratio between the slopes is close to constant with depth, which illustrates that $\Delta U\sim\exp(2k_{0}z)$ near the surface in these cases. The scaling is not perfect, particularly at the shallowest depths; this should not be surprising since the depth dependence of wave–turbulence interaction should scale not only with the wavelength, but also the turbulent integral length scale, which delimits the vertical extent of the topmost layer where the kinematic boundary condition at the surface begins to limit the vertical extent of turbulent eddies (the blocking effect, see, e.g., Teixeira & Belcher, 2002). Second, while the value of $\Delta U(z)$ after the passing of a group depends strongly on the turbulence level and steepness, there is no such trend for the relative slope ($\Delta U^{\prime}(z)/u_{s}^{\prime}(z)$), Case 1.A excepted.
@@
In Section 4.4 we will develop a RDT model describing the early onset of wave–turbulence interaction, which we can compare with the measurements in figure 3b, given this evidence that the combined wave/turbulence flow is still far from fully developed in Experiment 1.
@@
In conclusion, the evidence suggests that the change in Eulerian-mean current observed in our experiments is due to wave–turbulence interaction, and increases with increasing turbulence and increasing wave steepness.
@@ 3.3 Velocity change under regular waves (Experiments 2 and 3)
Cases 2.A3.B all consider turbulence interacting with regular (i.e., continuous and periodic) waves. For cases 2.A.1–2.B.2, we evaluate the mean streamwise velocity in the presence of waves, and subtract off the mean velocity profile from the ambient flow case without waves. This velocity difference we define as $\Delta U(z)$. Although the flow is wavy, the time series is long enough for the Eulerian-mean current, averaged over all $2000$ PIV images, to be well converged in the sense that further measurements would affect it insignificantly. See details in Appendix B.
@@
The turbulent current is affected by the waves during the time it takes it to traverse the test section, a distance $L_{\mathrm{FOV}}$ as defined in Sections 2.1.2 and 2.1.3. The interaction time is thus $T_{\mathrm{int}}=L_{\mathrm{FOV}}/U_{0}$, approximately $29$ s for Experiment 2 and $46$ s for Experiment 3, the latter having a slower flow speed. This corresponds to a number of wave cycles, $T_{\mathrm{int}}/T_{0},$ between $36$ (Cases 2.A and 3.A) and $83$ listed in table 4. Due to the slower flow speed. Cases 3.A and 3.B interact for considerably longer than 2.A and 2.B, both in terms of absolute time (in seconds) and in terms of number of intrinsic wave periods $T_{0}$, i.e., $T_{\mathrm{int}}/T_{0}$ is the number of wave cycles that the turbulence has encountered.
@@
Figure 4: The wave-induced current $\Delta U$ under regular waves as a function of depth $k_{0}z$ for cases 2.A.1, 2.A.2, 2.B.1, 2.B.2, 3.A and 3.B in panels (a)–(f), respectively. For each case, different wave steepness values $ak_{0}$ are shown as indicated in the legend. The dashed lines are the theoretical Stokes drift profiles at the same location for each case, shown as $u_{s}(z)$, that is, with opposite sign to the Stokes drift. The filled circles at $k_{0}z=-4$ and 0 indicate the theoretical value of the Eulerian return flow, $u_{\mathrm{rf}}$.
@@
In figure 4 we show $\Delta U(z)$ for cases 2.A3.B, where the different wave steepness values $ak_{0}$ are labelled in the legend. Several characteristic behaviours can be observed. First, all graphs have similar dependence on depth, with the highest absolute value nearest to the surface, decreasing down to a level where $k_{0}z$ is roughly in the range between 1-1 and 2-2, then turning and slowly becoming more negative again. (The non-monotonicity (turning) would not be visible in figure 3, where measurements were only made for $k_{0}z>-1.2$). Indeed, some cases see the induced Eulerian-mean current take positive values over a certain depth range. (One may note that the current profile is similar in shape and magnitude to those of Rashidi et al., 1992). Second, there is again a clear tendency that higher steepness leads to a larger vertical variation of $\Delta U(z)$. Finally, there appears to be a near-constant offset between graphs (an addition to the mean flow), which increases with steepness, and we will find in Section 4.1 that it can be explained, at least in part, as the approximately depth-uniform Eulerian return flow $u_{\mathrm{rf}}$.
@@
The reverse of the theoretical Stokes drift $u_{s}(z)$ from (1) is plotted in all panels of figure 4 with dashed lines of the same colour for comparison, calculated for each case. For all cases we note that $|\Delta U|<u_{s}$ nearest the surface, but that the depth variation $\mathrm{d}|\Delta U|/\mathrm{d}z\sim\mathrm{d}u_{s}/\mathrm{d}z$, is higher in some cases, smaller in others. We will return to these points later when comparing the results with theory.
@@
To further illustrate how the Eulerian-mean-current change depends on wave steepness, we plot the value of $\Delta U$ at a set reference depth as a function of steepness $ak_{0}$ in figure 5. The value at the shallowest depth available for all cases is used, i.e. $k_{0}z=-0.27$. For cases 2.A.1 and 2.B.1, $\Delta U$ appears to scale as $(ak_{0})^{2}$, indicated by the dashed line. Cases 2.A.2 and 2.B.2 do not adhere to the scaling, possibly due to the high wave steepness values involved, so that interactions of order $(ak_{0})^{3}$ and higher may become significant. A further possible explanation is the lack of scale separation: the wavelength is likely only slightly larger than the integral length scale in these two cases (based on comparison with cases 1.A1.D in table 2, which have similar $U_{0}$ – unfortunately, a direct comparison of $L_{x}^{x}$ is not possible, as explained). The comparatively large integral length scale and high turbulence levels mean that stronger angular scattering of waves on turbulent velocity changes is to be expected (Villas Bôas & Young, 2020; Smeltzer et al., 2023), a process which does not scale with steepness. Figure 5 should be interpreted only qualitatively, since the value of $\Delta U$ at a constant value of $k_{0}z$ is not entirely comparable between cases with different turbulence properties. As discussed in connection with figure 3(c) and at length in Section 4.4, $\Delta U$ depends not only on $k_{0}$ but also on the turbulent integral length scale and anisotropy, which varies between cases.
@@
Figure 5: The wave-induced current under regular waves at the ‘reference’ depth $k_{0}z=-0.27$ as a function of wave steepness for cases 2.A3.B as indicated in the legend. The dashed line is proportional to $(ak_{0})^{2}$.